Susan rolls a number cube with the numbers 1-6 on the faces.What is the probability that the number rolled is NOT even?
Answer by jim_thompson5910(35256)

Note: there are 3 odd numbers from 1-6 which are 1, 3, and 5

P(NOT even) = P(odd) = 3/6 = 1/2

So the probability that the number rolled is NOT even is 1/2 (which is 0.5 or 50%)